# InfoAffect: A Dataset for Affective Analysis of Infographics

### v1

@@ 4.2.1 Study Procedure
  • Do you think this kind of affective feedback could be useful for future design processes?
Interviews were conducted with participants in a quiet setting, audio-recorded with informed consent, and supplemented by contemporaneous field notes capturing key quotations, clarifications, and contextual observations. Each session lasted on average about 25 minutes per participant. All recordings and notes were subsequently transcribed and anonymized for analysis. Our analysis focused on how participants interpreted the detected affects, the reasoning they offered for those interpretations, and the extent to which these insights could inform affect-aware design practices. We also looked for recurring patterns and contrasts across roles to surface implications for future tool and workflow design.
4.2.2 Intent vs. Observation
We contrasted each participant’s original intent for their infographic with the affect extracted by the InfoAffect dataset construction pipeline. We then discussed these contrasts in semi-structured interviews. Representative examples are shown in Fig. 5. The contrasts fall into two patterns:
Validating
[…]
Correcting
The correction aspect focuses on situations where discrepancies arise between the creator’s intended affect and the extracted affects. Instead of directly modifying the design, the dataset provides an affective reference that helps creators recognize potential inconsistencies between their intended and extracted affects. Participants are then encouraged to adjust their own works accordingly—either by reinforcing insufficient affects or by reducing unintended affects.
[…]
4.3 Exp. II : Evaluating accuracy of InfoAffect dataset
This section introduced the Exp.II, which aimed to quantitatively assess the accuracy of affective extraction performed by InfoAffect dataset. The goal was to verify whether the affects identified by the system were consistent with human perception when evaluated by a general audience.
4.3.1 Study Procedure
Two sources of infographics were selected for this study: ❶ 430 items randomly sampled from the constructed InfoAffect dataset. ❷ 10 infographics collected in Exp I, in which discrepancies were observed between the participants’ intended affects and the results of InfoAffect dataset construction pipeline. In total, 305 infographics were included in the study. We used Amazon Mechanical Turk (MTurk) to gather responses from a large and diverse pool of participants. To mitigate variability from English proficiency, we limited eligibility to U.S.-based MTurk workers with a $\geq 90\%$ approval rate on prior tasks. 32 participants were permit to complete the affective evaluation task after a 5-minute standardized introduction. Participation was voluntary and we paid each participant \$1.50 for labeling 10 infographics. Responses were analyzed in aggregate only. Additionally, we included the 12 participants from Exp I, yielding a total sample of 44 unique participants. Each infographic was presented to participants for affective evaluation.
Since the second-level classification produceed a large number of fine-grained affects, it was not suitable for direct use in a three-option affect selector. Therefore, we adopted the first-level classification results as the basis for evaluation and used a three-option affect selector: $\text{}\square$ Negative, $\text{}\square$ Neutral, $\text{}\square$ Positive for each infographic. This design allows each infographic to be independently evaluated on the intensity of positive, neutral, and negative affects. Each infographic received ratings from multiple annotators. The accuracy was determined by comparing whether the human judgments matched the corresponding affects assigned in InfoAffect dataset.
4.3.2 CACI Analysis
To evaluate the agreement between the system-detected and human-rated affective results, we proposed a Composite Affect Consistency Index (CACI) that combines the dominant affect agreement and the overall distribution similarity. It was defined as:
$CACI=M\times S$ (6)
To prepare the comparison, we represent the results as three-dimensional vectors for the Negative, Neutral, and Positive classes, where each entry records the number of raters selecting that class
$\mathbf{v}_{1}=(n_{1},ne_{1},p_{1}),\qquad\mathbf{v}_{2}=(n_{2},ne_{2},p_{2}).$
Given these vectors, we let the dominant class be the index of the largest entry in each vector and define
$M=\begin{cases}1,&\text{if }\operatorname*{arg\,max}(n_{1},ne_{1},p_{1})=\operatorname*{arg\,max}(n_{2},ne_{2},p_{2}),\\ 0,&\text{otherwise.}\end{cases}$ (7)
We compute $S\in[0,1]$ as the cosine similarity between $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$:
$S=\frac{n_{1}n_{2}+ne_{1}ne_{2}+p_{1}p_{2}}{\sqrt{n_{1}^{2}+ne_{1}^{2}+p_{1}^{2}}\,\sqrt{n_{2}^{2}+ne_{2}^{2}+p_{2}^{2}}}.$ (8)
A higher value of CACI indicates stronger consistency, both in terms of the dominant affect and the overall affective tendency. In cases where the dominant affect differs ($M=0$), the index naturally dropped to zero, emphasizing disagreement in primary affect.
As shown in Fig. 5, we presented infographic samples together with their corresponding CACI values. In general, most samples exhibited high CACI values, indicating that the InfoAffect dataset was largely consistent with the crowd-sourced ratings. Then, we computed the mean CACI values within the Negative, Neutral, and Positive classes. The results showed that the Positive class achieved the highest consistency ($\overline{CACI}_{pos}=0.992$), followed by the Neutral class ($\overline{CACI}_{neu}=0.982$), while the Negative class exhibited slightly lower agreement ($\overline{CACI}_{neg}=0.985$). Among all evaluated samples, 68% of the samples scored above 0.9, while only 7% fell below 0.8. Finally, we aggregated all samples to compute the CACI of InfoAffect dataset. The resulting $\overline{CACI}_{InfoAffect}=0.986$. These results collectively demonstrated that the affective recognition maintains high accuracy in identifying affects of infographics.
5 Limitations and Future Work
While InfoAffect dataset advanced affective analysis of infographics, we acknowledged two important limitations. First, our dataset size remains modest, which limits coverage across topics, styles, and rare affect categories. This limitation place natural bounds on generalizability and reduces ability to characterize subtle edge cases. Second, while the first-level classification results are balanced, the second-level classification results are long-tailed due to the fine-grained classification and the use of real-world infographics. Such imbalance complicates training procedures and evaluation, often increasing variance. These issues are typical for real-word data and they underscore that our results should be read with appropriate caution, especially for low-frequency affects.
Element-wise affect controls.
We will study how specific visual and textual elements in infographics shape perceived affect by varying one factor at a time on existing designs and noting the affect shifts. The focus is on color palettes, type choices, icon shapes, layout density, chart encodings, headlines, and captions. We will record patterns that hold across examples and note when elements interact. The outcome will be a brief set of practical guidelines and small before–after examples showing what to change when a designer seeks caution, neutrality, or a mild positive tone, with a simple format suitable for everyday editing and classroom use.
Standardized evaluation and scalable data.
We will develop a simple, shared methodology for measuring affects in infographics, such as a small set of rating questions and scales, clear rater instructions with controlled viewing time, anchor stimuli for calibration, and a short reporting template that records who rated, what was shown, exposure duration, and how scores were computed. We will release this as a lightweight kit to improve comparability, reuse, and replication across labs and projects. In parallel, we will curate a larger-scale, more diverse dataset for fine-tuning multimodal large language models—covering broader topics, languages, and visual styles, with better balance for rare affects and element-level tags plus brief rater notes to clarify labels.
6 Conclusion
We presented InfoAffect dataset, a 3.5k-sample dual-modal that pairs real-world infographics with accompanied text and constrains affect labels with the Affect Table. The corpus covers six topical fields and is prepared with three methods to ensure alignment and quality. We extract affects with five state-of-the-art MLLMs and fuse their outputs using Reciprocal Rank Fusion to obtain robust affective annotations. We validated the resource with two experiments that assessed usability and accuracy of InfoAffect dataset. The accuracy study introduced the Composite Affect Consistency Index and achieved an overall score of 0.986. InfoAffect dataset establishes a foundation for future research on affective understanding, model training, and affectively informed visualization design.
Acknowledgements.

### v2

@@ Appendix F Interview transcripts
Q5: Do you think this kind of affective feedback could be useful for future design processes?
A5: Yes, very much so. Such feedback would help refine the affective balance in policy communication, ensuring that data visualizations motivate sustainable action without overstating superiority or competitiveness.
eg.2 E2
<original intent>: I created this infographic to objectively present global viewing preferences for horror movies and series, illustrating which countries show the highest engagement with this genre. My intent was primarily informative rather than affective, aiming to highlight a cultural trend through clean visual comparison.
<extracted affects>: Fearful, intense.
Q1: Do you think the affect identified aligns with the affect you originally intended to convey in this infographic?
A1: Not exactly. The system saw it as “sinister” or “fearful” but I only meant to report viewing data with subtle horror-themed coherence, not create an unsettling vibe.
Q2: If the extracted affects differs from your original intent, how would you describe that difference?
A2: The system exaggerated the tone. I wanted mild thematic links to horror, not to directly evoke fear as it interpreted.
Q3: Do you think the visual elements, such as color, composition, or typography, contributed to this unintended affects?
A3: Probably. The dark background, bright red bars, and red balloon likely amplified fear associations, making it more disturbing than intended.
Q4: Does the analysis results help you better understand how affect is represented in your visual design?
A4: Yes. Even small stylistic choices can strongly shape affective interpretation, despite neutral data.
Q5: Do you think this kind of affective feedback could be useful for future design processes?
A5: Absolutely. It’ll help me balance “genre flavor” and avoid unintended affective responses while keeping thematic relevance.
eg.3 P2
<original intent>: I designed this infographic to show global differences in reported mental health symptoms in a clear and empathetic way. My main goal was to raise awareness, not to dramatize the data.
<extracted affects>: Calm, empathetic, serious.
Q1: Do you think the affect identified aligns with the affect you originally intended to convey in this infographic?
A1: Yes, completely. The system picked up on exactly what I wanted, a calm and empathetic tone that shows the data respectfully.
Q2: If the extracted affects differs from your original intent, how would you describe that difference?
A2: [The extracted affect aligns with the original intent.]
Q3: Do you think the visual elements, such as color, composition, or typography, contributed to this unintended affects?
A3: [The extracted affect aligns with the original intent.]
Q4: Does the analysis results help you better understand how affect is represented in your visual design?
A4: Yes, it confirms that even simple design choices, like color and icon style, can really shape how people affectively read the data.
Q5: Do you think this kind of affective feedback could be usefulfor future design processes?
A5: For sure. Getting this kind of feedback helps me make sure I’m hitting the right affective tone, especially for sensitive topics like mental health.
eg.4 P3[Uncaptioned image] <original intent>: I made this infographic simply to list the relevant figures in my manuscript, keeping the tone descriptive rather than evaluative. <extracted affects>: alarming, concerning.
Q1: Do you think the affect identified aligns with the affect you originally intended to convey in this infographic?
A1: No. My original intent was to keep the tone descriptive rather than evaluative, but the identified affect was negative.
Q2: If the extracted affects differs from your original intent, how would you describe that difference?
A2: I intended to simply list the relevant figures, but the detected affect appeared “alarming” and “concerning” which differs from her neutral, factual intent.
Q3: Do you think the visual elements, such as color, composition, or typography, contributed to this unintended affects?
A3: Yes. I reflect that it’s due to the warm red–orange ramp and the strong contrast in the legend.
Q4: Does the analysis results help you better understand how affect is represented in your visual design?
A4: Yes. It helps me realize how color and contrast can unintentionally convey negative affectives even in descriptive visualizations.
Q5: Do you think this kind of affective feedback could be usefulfor future design processes?
A5: Yes. Such feedback could help designers anticipate and adjust unintended affective tones in their visual work.
eg.5 P4
<original intent>: I submitted this infographic aimed to naturally present the information, simply enumerated which countries are more open to electric vehicles without taking an evaluative stance.
<extracted affects>: informational.
Q1: Do you think the affect identified aligns with the affect you originally intended to convey in this infographic?
A1: Yes, perfectly. My goal was just to report which nations are considering electric vehicles, the analysis produced neutral terms that align with the report-style tone I intended.
Q2: If the extracted affects differs from your original intent, how would you describe that difference?
A2: [The extracted affect aligns with the original intent.]
Q3: Do you think the visual elements, such as color, composition, or typography, contributed to this unintended affects?
A3: [The extracted affect aligns with the original intent.]
Q4: Does the analysis results help you better understand how affect is represented in your visual design?
A4: Yes. It confirms that using simple bar graphs, neutral colors, and clear labels can effectively convey an informational tone, helping me grasp how visual choices shape affect representation.
Q5: Do you think this kind of affective feedback could be usefulfor future design processes?
A5: For sure. It ensures I maintain the neutral, informative tone I want.
eg.6 P5
<original intent>: I made this infographic to simply report the world’s busiest airports’ passenger traffic data, aiming for a neutral, factual tone.
<extracted affects>: Informational, neutral.
Q1: Do you think the affect identified aligns with the affect you originally intended to convey in this infographic?
A1: Yes, totally. I just wanted to lay out the airport traffic facts, and the system picked up “informational, neutral” vibes, exactly what I intended.
Q2: If the extracted affects differs from your original intent, how would you describe that difference?
A2: [The extracted affect aligns with the original intent.]
Q3: Do you think the visual elements, such as color, composition, or typography, contributed to this unintended affects?
A3: [The extracted affect aligns with the original intent.]
Q4: Does the analysis results help you better understand how affect is represented in your visual design?
A4: Definitely. It shows that using straightforward visuals and a neutral color scheme can nail an informational tone, so I get how design choices shape that now.
Q5: Do you think this kind of affective feedback could be usefulfor future design processes?
A5: For sure. It helps me make sure I keep that neutral, data-driven tone in future infographics where I just want to present facts.
eg.7 P6
<original intent>: I offered this infographic aiming for “hopeful”, seeking a positive affect.
<extracted affects>: optimistic.
Q1: Do you think the affect identified aligns with the affect you originally intended to convey in this infographic?
A1: Yes. My intent in the policy report was to highlight a hopeful trend for policymakers’ consideration, and the extracted affect of the infographic was “optimistic” which closely aligned with my original intent.
Q2: If the extracted affects differs from your original intent, how would you describe that difference?
A2: [The extracted affect aligns with the original intent.]
Q3: Do you think the visual elements, such as color, composition, or typography, contributed to this unintended affects?
A3: [The extracted affect aligns with the original intent.]
Q4: Does the analysis results help you better understand how affect is represented in your visual design?
A4: Definitely. It confirms that using upward rending lines and distinct colors can effectively convey an optimistic, hopeful tone, deepening my understanding of how visual elements shape affect representation.
Q5: Do you think this kind of affective feedback could be usefulfor future design processes?
A5: For sure. It helps me ensure that the optimistic, hopeful tone I intend is consistently conveyed in future designs, especially when highlighting positive trends like this one.
eg.8 P7
<original intent>: I created this infographic to show the global situation of people finding healthy food too expensive, aiming to present the data in a somewhat approachable and relatable way.
<extracted affects>: Informative, concerned.
Q1: Do you think the affect identified aligns with the affect you originally intended to convey in this infographic?
A1: Not really. I wanted it to be more approachable, but the system picked up an informative and concerned tone.
Q2: If the extracted affects differs from your original intent, how would you describe that difference?
A2: My intent was to make it relatable, but the system’s identified affect is more focused on being informative and concerned, missing the intended approachable vibe.
Q3: Do you think the visual elements, such as color, composition, or typography, contributed to this unintended affects?
A3: Yeah. The orange bar graph, country flags, and the serious “\$\$\$” and broccoli icon make it feel informative and concerned, which is not as approachable as he wanted.
Q4: Does the analysis results help you better understand how affect is represented in your visual design?
A4: Definitely. It shows that even if you want an approachable tone, the visual choices like bar style and icons can push it toward a more informative and concerned affect.
eg.9 P10
<original intent>: I made this infographic to express a negative winner-takes-all trend, namely a worrying concentration in which one platform dominates the others.
<extracted affects>: confident, competitive.
Q1: Do you think the affect identified aligns with the affect you originally intended to convey in this infographic?
A1: No. My intent is to express a negative winner-takes-all trend, but the identified affect was positive.
Q2: If the extracted affects differs from your original intent, how would you describe that difference?
A2: I want to show “a worrying concentration in which one platform dominates” yet the extracted affects are “confident” and “competitive” giving a sense of achievement rather than imbalance.
Q3: Do you think the visual elements, such as color, composition, or typography, contributed to this unintended affects?
A3: Yes. I reflected that the bright, saturated palette and clean ranking layout make it feel like achievement rather than imbalance, and the long red top bar reads as “leading” more than “concerning”, while the crisp flags and labels add a celebratory tone.
Q4: Does the analysis results help you better understand how affect is represented in your visual design?
A4: Yes. It reveals how design choices such as palette, layout, and iconography can shift the affective reading from critical to celebratory.
Q5: Do you think this kind of affective feedback could be usefulfor future design processes?
A5: For sure. Such feedback can help designers better align visual tone with their intended message and avoid unintended positive framing of negative topics.
Q5: Do you think this kind of affective feedback could be usefulfor future design processes?
A5: For sure. It helps me see where his design choices don’t match my intended tone, so he can adjust things like visuals or color to get the affect he wants next time.